OpenStudy (anonymous):

What are the x-intercept(s) of the graph of: y − x2 = −11x + 24? (−3, 0) and (8, 0) (3, 0) and (−8, 0) (−3, 0) and (−8, 0) (3, 0) and (8, 0)

OpenStudy (anonymous):

a brief explanation would b nice too :)

OpenStudy (cwrw238):

the x intercepts are the values of x when y = 0 so you need to solve the equation x^2 - 11x + 24 = 0

OpenStudy (anonymous):

solve for x?

OpenStudy (cwrw238):

can you solve this?

OpenStudy (anonymous):

i just dont understand what u want me to solve? x?

OpenStudy (cwrw238):

yes - solve for x - have you done this type of problem before?

OpenStudy (cwrw238):

yes - solve for x - have you done this type of problem before?

OpenStudy (anonymous):

no thats y im confused

OpenStudy (cwrw238):

ok - all another way to do this problem is to plug in the given coordinates and see which ones fit

OpenStudy (anonymous):

This the sort of problem you'd have to graph, so looking for the point or the point slop-intercept is your objective. Start with each coordinate and plug them into the equation to see which makes the most sense, that'll give you your answer...tell me if that works for you :D lol

OpenStudy (anonymous):

@brookester6 can u gimme an example of how to plug the coordinates in?

OpenStudy (cwrw238):

e.g. try (-3,0 ) and (8 , 0 ) that is when x = -3 or x=8 y = 0 (-3)^2 - 11(-3) + 24 = 9 + 33 + 24 = 66 this should be 0 doesn't fit so we needn't try x = 8

OpenStudy (anonymous):

then we try (3,0) and (-8,0)? right

OpenStudy (anonymous):

quick question where did u get the =9+33+24=66 from?

OpenStudy (anonymous):

*@cwrw238 quick

OpenStudy (anonymous):

For example...y − x2 = −11x + 24?...and the coordinates you want to plug in are (3,0) and (-8,0) so instead of writing an equation, you would use this...m=x-x2/y-y1

OpenStudy (anonymous):

so then use those coordinates and plug them in

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

np :)